# Flexibility Pricing in Distribution Systems: A Direct Method Aligned with Flexibility Models

### v1

@@ V-C Impact of the Aggregators’ Flexibility Costs on the Market Clearing Results
Next, we test the case where only the bidding cost coefficients of one aggregator change. It is assumed that the aggregators’ original cost coefficients are their true cost coefficients. Keeping the cost coefficients of other aggregators constant, we change the bidding cost coefficients of the aggregator at node 2 by multiplying it with the parameter $\beta$, implying that this aggregator does not bid its true cost coefficients. After the DSO flexibility market is cleared, the net profit of the aggregator (the received payment of the aggregator $R_{h}$ minus the aggregator’s true flexibility cost) can be obtained, as shown in Fig. 9. It is observed that the maximum net profit for the aggregator occurs at $\beta=1$, i.e., when the aggregator bids its true flexibility costs. This result verifies that the proposed market mechanism can encourage aggregators to report their true cost coefficients in a sufficiently competitive environment.
Figure 9: Changes in the profit of the aggregator at node 2 with its bidding flexibility cost.
@@ VI Conclusions
This paper proposes a flexibility pricing method aligned with the flexibility models of DERs. For the flexibility of individual DERs, we propose to formulate its flexibility cost based on adjustment ranges in both power and accumulated energy consumption, which extends the traditional power range-based formulations. The proposed flexibility costs formulation quantifies the potential inconvenience caused to users of typical DERs such as EVs, DESS, and HPs by flexibility activation. We also propose a flexibility cost formulation aligned with the aggregated flexibility model and a DSO flexibility market model for activating flexibility to participate in the transmission system operation for energy arbitrage and reserve capacity provision. Furthermore, we propose the concept of MFPs for settlement, which ensures incentive compatibility, revenue adequacy, and a non-profit DSO. Numerical results validate the proposed pricing and market-clearing methods.
A potential limitation of this work lies in the calculation of the aggregators’ cost coefficients, which are estimated using a weighted average method in this paper and may be improved in future works. Future works can also focus on the coordination of the transmission system operation and the DSO flexibility market to better utilize flexibility to improve the operational efficiency of the entire power system.
Appendix A Proof of Proposition 1
We first reformulate the temperature evolution model (5) using a time-sequential form, as follows:
@@ Appendix A Proof of Proposition 1
$\boldsymbol{\theta}={\mathbf{A}}{\boldsymbol{\theta}^{{\text{amb}}}}+\frac{\eta}{H}\cdot{\mathbf{Ap}}+{\boldsymbol{\alpha}}{\theta_{0}},$ (A.2)
@@
$\boldsymbol{\theta}^{\text{set}}={\mathbf{A}}{\boldsymbol{\theta}^{{\text{amb}}}}+\frac{\eta}{H}\cdot{\mathbf{Ap}^{\text{base}}}+{\boldsymbol{\alpha}}{\theta_{0}}.$ (A.3)
@@
$\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}=\frac{\eta}{H}\cdot{\mathbf{A}(\mathbf{p}-\mathbf{p}^{\text{base}})}.$ (A.4)
@@
${\mathbf{p}}-\mathbf{p}^{\text{base}}=\frac{H}{\eta}{{\mathbf{A}}^{-1}}\left({{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\text{set}}}\right).$ (A.5)
@@
$-\Delta\hat{\boldsymbol{\theta}}\leq\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}\leq\Delta\check{\boldsymbol{\theta}}.$ (A.6)
This constraint cannot be transformed into power adjustment ranges in each time slot because the components of $\mathbf{A}^{-1}$ are not all positive, rendering the operation of multiplying both sides of inequality (A.6) by $\mathbf{A}^{-1}$ invalid. [^3] Applying the Fourier-Motzkin Elimination to (A.5) and (A.6) can indeed derive the constraints of $\mathbf{p}-\mathbf{p}^{\text{base}}$, but the resulting constraints do not form upper and lower bounds for $\mathbf{p}-\mathbf{p}^{\text{base}}$ in each time slot. Therefore, we aim to establish a connection between the energy adjustment ranges $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$ and the temperature adjustment ranges $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$.
To this end, we first write the matrix form of the energy-power relationship (4):
@@
${\mathbf{e}}-\mathbf{e}^{\text{base}}=\Delta_{T}\cdot\frac{H}{\eta}{{\mathbf{LA}}^{-1}}\left({{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\text{set}}}\right),$ (A.8)
@@
All the components of ${\mathbf{L}}{{\mathbf{A}}^{-1}}$ are positive since $10<a<$ 1 by definition, hence it is feasible to multiply both sides of inequality (A.6) by $\Delta_{T}\cdot\frac{H}{\eta}{\mathbf{L}}{{\mathbf{A}}^{-1}}$ to derive the constraints on the energy adjustment ${\mathbf{e}}-\mathbf{e}^{\text{base}}$, i.e.,
$-\Delta_{T}\cdot\frac{H}{\eta}{\mathbf{L}}{{\mathbf{A}}^{-1}}\Delta\hat{\boldsymbol{\theta}}\leq{\mathbf{e}}-\mathbf{e}^{\text{base}}\leq\Delta_{T}\cdot\frac{H}{\eta}{\mathbf{L}}{{\mathbf{A}}^{-1}}\Delta\check{\boldsymbol{\theta}}.$ (A.9)
Let $\mathbf{D}\triangleq\Delta_{T}\cdot\frac{H}{\eta}{\mathbf{L}}{{\mathbf{A}}^{-1}}$, then $\Delta\hat{\mathbf{e}}={\mathbf{D}}\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\mathbf{e}}={\mathbf{D}}\Delta\check{\boldsymbol{\theta}}$, which complete the proof of Proposition 1.

### v4

@@ V-C Impact of the Aggregators’ Flexibility Costs on the Results
Next, we test the case where only the reported cost coefficients of one aggregator change. It is assumed that the aggregators’ original cost coefficients are their true cost coefficients. Keeping the cost coefficients of other aggregators constant, we change the reported cost coefficients of the aggregator at node 2 by multiplying it with the parameter $\beta$, implying that this aggregator does not report its true cost coefficients. After the DSO optimization problem is solved, the net profit of the aggregator (the received payment of the aggregator $R_{h}$ minus the aggregator’s true flexibility cost) can be obtained, as shown in Fig. 9. It is observed that the maximum net profit for the aggregator occurs at $\beta=1$, i.e., when the aggregator reports its true flexibility costs. This result verifies that the proposed flexibility activation model can encourage aggregators to report their true cost coefficients.
Figure 9: Changes in the profit of the aggregator at node 2 with its reported flexibility cost.
@@ VI Conclusions
This paper proposes a quantification method for the costs of the time-coupled flexibilities from DERs. For individual DERs, we propose to compute their flexibility cost based on adjustment ranges in both power and accumulated energy consumption. From the aggregator’s standpoint, we propose to compute their flexibility cost based on adjustment ranges in the flexibility trajectory. The proposed quantification method is unified for individual and aggregated DERs. It can also capture the time-coupled characteristics of the flexibility regions, which are not reflected in the traditional power-range-based approaches.
A DSO optimization model is then designed to activate the flexibility to participate in the transmission system operation for energy arbitrage and reserve capacity provision. From this optimization model, we derive the concept of MFPs, which quantify the marginal contributions of the activated flexibility to the DSO’s total revenue in the TSO-DSO coordination (the value of flexibility). The MFPs are then used to allocate the revenues obtained from the TSO to each aggregator in the distribution system. The proposed DSO optimization model and the revenue allocation strategy can ensure incentive compatibility, revenue adequacy, and a non-profit DSO.
A potential limitation of this work lies in the calculation of the aggregators’ cost coefficients, which are estimated using a weighted average method in this paper and may be improved in future works. Future work can also focus on real-time disaggregation and control to better exploit flexibility in TSO-DSO coordination. [Proof of Proposition 1] We first reformulate the temperature evolution model (5) using a time-sequential form, as follows:
@@
$\boldsymbol{\theta}=(1-\alpha){\mathbf{A}}{\boldsymbol{\theta}^{{\text{amb}}}}+(1-\alpha)\frac{\eta}{H}\cdot{\mathbf{Ap}}+{\boldsymbol{\alpha}}{\theta_{0}},$ (A.2)
@@
$\boldsymbol{\theta}^{\text{set}}=(1-\alpha){\mathbf{A}}{\boldsymbol{\theta}^{{\text{amb}}}}+(1-\alpha)\frac{\eta}{H}\cdot{\mathbf{Ap}^{\text{base}}}+{\boldsymbol{\alpha}}{\theta_{0}}.$ (A.3)
@@
$\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}=(1-\alpha)\frac{\eta}{H}\cdot{\mathbf{A}(\mathbf{p}-\mathbf{p}^{\text{base}})}.$ (A.4)
@@
${\mathbf{p}}-\mathbf{p}^{\text{base}}=\frac{H}{\eta(1-\alpha)}{{\mathbf{A}}^{-1}}\left({{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\text{set}}}\right).$ (A.5)
@@
$-\Delta\check{\boldsymbol{\theta}}\leq\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}\leq\Delta\hat{\boldsymbol{\theta}}.$ (A.6)
We first prove the second half of Proposition 1 using a contradiction. Suppose there exists an $\mathbf{F}\in\mathbb{R}^{T\times T}$ that satisfies the conditions. We set $\Delta\hat{\boldsymbol{\theta}}=\mathbf{0}$ and $\Delta\check{\boldsymbol{\theta}}>\mathbf{0}$, so that the power profile $\mathbf{p}$ should satisfy $-\mathbf{F}\Delta\check{\boldsymbol{\theta}}\leq\mathbf{p}-\mathbf{p}^{\text{base}}\leq\mathbf{0}$. Let $\theta_{1}=\theta_{1}^{\text{set}}-\Delta\check{\theta}_{1}$ and $\theta_{2}=\theta_{2}^{\text{set}}$, then the first two rows of Equation (A.5) becomes:
$\left[{\begin{array}[]{cc}{{p_{1}}-p_{1}^{{\text{base}}}}\\ {{p_{2}}-p_{2}^{{\text{base}}}}\end{array}}\right]$ $=\frac{H}{{\eta(1-\alpha)}}\left[{\begin{array}[]{cc}1&{}\hfil\\ {-\alpha}&1\end{array}}\right]\left[{\begin{array}[]{cc}{-\Delta{\theta_{1}}}\\ 0\end{array}}\right]$ (A.13)
$=\frac{H}{{\eta(1-\alpha)}}\left[{\begin{array}[]{cc}{-\Delta{\check{\theta}_{1}}}\\ {\alpha\Delta{\check{\theta}_{1}}}\end{array}}\right].$ (A.16)
Thus, ${{p_{2}}-p_{2}^{{\text{base}}}}=\frac{H\alpha\Delta{\check{\theta}_{1}}}{{\eta(1-\alpha)}}>0$, which contradicts $\mathbf{p}-\mathbf{p}^{\text{base}}\leq\mathbf{0}$, proving the second half of Proposition 1.
Readers might wonder why we cannot multiply both sides of (A.6) by $\frac{H}{\eta(1-\alpha)}{{\mathbf{A}}^{-1}}$ to generate a constraint on $\mathbf{p}-\mathbf{p}^{\text{base}}$. This is because the elements of ${\mathbf{A}}^{-1}$ are not all positive, and thus, multiplying both sides of (A.6) by $\frac{H}{\eta(1-\alpha)}{{\mathbf{A}}^{-1}}$ does not necessarily preserve the inequality’s direction. Nevertheless, we will show that this operation becomes feasible when transforming the temperature adjustment ranges $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$ into the energy adjustment ranges $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$, and it will prove the first half of Proposition 1. To this end, we write the matrix form of the energy-power relationship (4):
@@
${\mathbf{e}}-\mathbf{e}^{\text{base}}=\frac{\Delta_{T}H}{\eta(1-\alpha)}{{\mathbf{LA}}^{-1}}\left({{\boldsymbol{\theta}}-\boldsymbol{\theta}^{\text{set}}}\right),$ (A.18)
@@
All the components of ${\mathbf{L}}{{\mathbf{A}}^{-1}}$ are non-negative since $10<a<$ 1 by definition, hence The direction of inequality (A.6) is preserved when multiplying both sides by $\frac{\Delta_{T}H}{\eta(1-\alpha)}{\mathbf{L}}{{\mathbf{A}}^{-1}}$, i.e.,
$-\frac{\Delta_{T}H}{\eta(1-\alpha)}{\mathbf{L}}{{\mathbf{A}}^{-1}}\Delta\hat{\boldsymbol{\theta}}\leq{\mathbf{e}}-\mathbf{e}^{\text{base}}\leq\frac{\Delta_{T}H}{\eta(1-\alpha)}{\mathbf{L}}{{\mathbf{A}}^{-1}}\Delta\check{\boldsymbol{\theta}}.$ (A.19)
Let $\mathbf{D}\triangleq\frac{\Delta_{T}H}{\eta(1-\alpha)}{\mathbf{L}}{{\mathbf{A}}^{-1}}$, then $\Delta\hat{\mathbf{e}}={\mathbf{D}}\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\mathbf{e}}={\mathbf{D}}\Delta\check{\boldsymbol{\theta}}$, which complete the proof of Proposition 1.